Is x=y^4+5 a function?
Answer by stanbon(75887)

Is x=y^4+5 a function?
--------
Solve for "y":

y = (x-5)^(1/4)
-----------------
Example: If you let x = 21 you get
y = (21-5)^(1/4)
y = 16^(1/4)
y = 2 or y = -2
---------------------
Since there are two y-values for a single value of "x"
x = y^4+5 is not a function.
====================================
Cheers,
Stan H.